Carla has $2.40 in nickels and dimes.Deron has $5.50 in dimes and quarters.
Deron has as many dimes as Carla has nickles and as many quarters as Carla has dimes/ How many of each coin does Carla have?

Answer by mananth(16946)

Carla has $2.40 in nickels and dimes.Deron has $5.50 in dimes and quarters.
Deron has as many dimes as Carla has nickles and as many quarters as Carla has dimes/ How many of each coin does Carla have?
..
let carla have x nickels & y dimes
5x+10y=240.................1
Deron has x dimes & y quarters
10x+25y=550.................2
...
multiply equation 1 by -2
-10x-20y=-480
add it to equation 2
10x-10x+25y-20y=550-480
5y=70
y=14 dimes carla has
plug value of y in equation 1
5x+140=240
5x=100
x=20 nickels carla has
.
m.ananth@hotmail.ca